French worth 9.00per pound, Kenyan worth 7.50 per pound want a 10 pound mixture worth 8.40 per pound how mauch of each would I need
1
Expert's answer
2011-05-17T09:32:54-0400
Let A be the total of French in the 10 pound mixture, then 10-A is the total of Kenyan.
Then the cost of this mixture is& 9A+7.5(10-A)& and should be equal to& 8.4*10=84.
So we have an equation:
9A+7.5(10-A)=84
9A-7.5A=84-75
1.5A=9
A=6.
Thus we have to take 6 pounds of French and 4 pounds of Kenyan